Find the general solution to X1 2 X1 -2 4 32 32
Answer #1

2 Uz с -> -( 1+j) 4,-24, 50 - (1+i) 4 = 242 let ₂= c then u = 2 (1+i) -2 (1-1) +1-1 = -2(1-1) = -(1-1) = (-1 + i) c. Hence. uthe solution that we get from the first eigenvalue and eigen vector is, (3+) H = 1 est (cost + isint) 7 X(t) = => X,(t) = /-e